OpenStudy (anonymous):

Use a half-angle formula to simplify: (sin4 theta)/(1+cos4theta).

OpenStudy (anonymous):

\[\sin4\theta/1+\cos4\]

zepdrix (zepdrix):

oh ok :) hmm

OpenStudy (anonymous):

The bottom half should read: (1+cos 4 theta)

zepdrix (zepdrix):

Ok so this appears to be the `Tangent Half-Angle Identity`. \[\large \tan \frac{x}{2}=\frac{\sin x}{1+\cos x}\]

zepdrix (zepdrix):

So we appears to have something that looks like the `right` side of this identity, yes?

zepdrix (zepdrix):

appear*

OpenStudy (anonymous):

right, yes.

zepdrix (zepdrix):

If it helps, maybe make a substitution. Let \(\large x=4\theta\) We know that this will simplify down to \(\large \tan \dfrac{x}{2}\). So from here, \(\large x=4\theta\), solve for x/2! :)

zepdrix (zepdrix):

is that confusing? :c

OpenStudy (anonymous):

Ok, never mind that post, I got it. Thanks!!

zepdrix (zepdrix):

k c: